A car slows from 23.69 m/s to rest in 4.44 s. How far did it travel in this time?


1
Expert's answer
2021-04-09T06:37:57-0400

Answer

Using

v=u+at

"0=23.69-a\times4.44\\a=5.35m/s^2"

So travelled distance

"S=\frac{v^2-u^2}{2a}\\=\frac{0-23.69^2}{-2\times5.35}\\=52.45m"